(b) If
A =


4 4 −2
−1 0 1
3 6 −1

 ,
show that A2 = A + kI for some constant k, where I is the unit matrix of order 3. Hence
find the inverse matrix A−1
. [
1
Expert's answer
2020-07-06T17:20:58-0400

Given, "A =\begin{bmatrix}
4 & 4 & -2 \\
-1 & 0 & 1\\
3 & 6 & -1
\end{bmatrix}".


"A^{2}=\begin{bmatrix}
4 & 4 & -2 \\
-1 & 0 & 1\\
3 & 6 & -1
\end{bmatrix} \cdot \begin{bmatrix}
4 & 4 & -2 \\
-1 & 0 & 1\\
3 & 6 & -1
\end{bmatrix}\\~\\~~~~~~ = \begin{bmatrix}
6 & 4 & -2 \\
-1 & 2 & 1\\
3 & 6 & 1
\end{bmatrix}\\~\\ ~~~~~~= \begin{bmatrix}
4 & 4 & -2 \\
-1 & 0 & 1\\
3 & 6 & -1
\end{bmatrix} + \begin{bmatrix}
2 & 0 & 0 \\
0 & 2 & 0\\
0 & 0 & 2
\end{bmatrix}\\~\\~~~~~~ = A + 2I."


Therefore, "A^{2}=A+kI" for "k=2".


Multiplying, "A^{2}=A+2I" by "A^{-1}", we get


"A = I + 2A^{-1}\\~\\
A^{-1} = \frac{1}{2}(A-I)\\~\\
~~~~~~~~=\dfrac{1}{2}\begin{bmatrix} 3 & 4 & 2\\ -1 & -1 & 1 \\3 & 6 & -2\end{bmatrix}"